OpenStudy (anonymous):

If you know the number of yards for a measurement, you can change that measure to meters by multiplying the number of yards by 0.9144. a. Write a function rule that relates meters to yards. b. How many meters are equivalent to 7,200 yards? Round your answer to the nearest hundredth, if necessary. c. How many yards are equivalent to 2,500 meters? Round your answer to the nearest hundredth, if necessary.

OpenStudy (anonymous):

Let's take 1 yard for this example. You multiply it by 0.9144, as the instructions say, to get a meter: \[(0.9144) \times 1 yard = 0.9144 meters\] So: 1 yard = 0.9144 meters Now we need to switch this around to find how many yards we get with one meter: \[(1 yard)/0.9144 = 1 meter\] Now let's make a function that returns yards (as f(x) ) when given meters (x): \[f(x) = (\frac{ 1 }{ 0.9144 }) x\] This function should be the one you need for a. (just plug in your amount of meters for x, and the answer is in yards). You can use it to solve for c. To solve for b., use this function: \[f(x) = 0.9144x\]